OpenStudy (anonymous):

Evaluate the definite integral: 4~1 5(sqrtx)dx

OpenStudy (anonymous):

i like how i just asked the question and you already got a medal for saying f5 haha

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

that notation would equate to an interval of [1,4] right?

OpenStudy (amistre64):

5 x^(1/2+2/2) 2.5.sqrt(x^3) ------------- = ------------ correct? 1/2 + 2/2 3

OpenStudy (anonymous):

\[\int\limits_{1}^{4}\]

OpenStudy (amistre64):

2.5.sqrt(x^3) ------------ ; x = 4 = 10.4.2/3 = 80/3 3

OpenStudy (amistre64):

10.xsqrt(x) ---------- ; x=1 = 10/3 3 80-10 ----- = 70/3 if i dint mess up lol 3

OpenStudy (anonymous):

lol you probably didn't, it was one of the answer choices =)

OpenStudy (anonymous):

what is f' of sin30degrees?, if that question makes sense

OpenStudy (amistre64):

f(x) = sin(30) you say? and find the derivative?

OpenStudy (amistre64):

sin(30) is a constant, so the derivative of a constant is 0

OpenStudy (amistre64):

by definition, constant remain unchanging ;) therefore their rate of change with respect to anything is 0

OpenStudy (anonymous):

Use differentials to estimate the value indicated. sin 27degrees i was trying to work it out and got 1/2 = f'(sin30degrees)*3degrees

OpenStudy (anonymous):

1/2+ ***